# DeDe: Detecting Backdoor Samples for SSL Encoders via Decoders

### v1

@@ Abstract
Self-supervised learning (SSL) is pervasively exploited in training high-quality upstream encoders with a large amount of unlabeled data. However, it is found to be susceptible to backdoor attacks merely via polluting a small portion of training data. The victim encoders mismatch triggered inputs with target embeddings, e.g., match the triggered cat input to an airplane embedding, such that the downstream tasks are affected to misbehave when the trigger is activated. Emerging backdoor attacks have shown great threats in different SSL paradigms such as contrastive learning and CLIP, while few research is devoted to defending against such attacks. Besides, the existing ones fall short in detecting advanced stealthy backdoors. To address the limitations, we propose a novel detection mechanism, DeDe, which detects the activation of the backdoor mapping with the cooccurrence of victim encoder and trigger inputs. Specifically, DeDe trains a decoder for the SSL encoder on an auxiliary dataset (can be out-of-distribution or even slightly poisoned), such that for any triggered input that misleads to the target embedding, the decoder outputs an image significantly different from the input. We empirically evaluate DeDe on both contrastive learning and CLIP models against various types of backdoor attacks, and demonstrate its superior performance over SOTA detection methods in both upstream detection performance and ability of preventing backdoors in downstream tasks.
@@ 1 Introduction
In this paper, we propose DeDe, an effective detector for backdoor samples in SSL settings. To address the limitations of existing detection methods and the emergence of increasingly advanced attacks, the proposed approach directly leverages the relationships between embeddings and images. DeDe aims to establish a region-to-region mapping to invert the normal mapping of an encoder, thereby enabling the identification of malicious behaviors that deviate from the normal mapping. Encoders’ normal utility can be considered as the correct mapping from similar data (usually from the same class) to a certain region of the embedding space [36]. Embeddings of similar images are observed to converge to a clustered region in the vector space; conversely, backdoor encoders aim to map triggered images to a different embedding region instead. Our approach exploits this property to find a bijective mapping between classes of images and classes of embeddings, where normal utility leads to intra-class mapping, while backdoor behaviors introduce inter-class mapping.
Overall, the proposed detection approach offers three significant advantages: 1) the proposed approach is effective and robust against various advanced backdoor attacks, which use the invisible trigger or stealthy embedding; 2) it does not rely on the availability of a clean dataset or an in-distribution dataset; and 3) it requires no prior information regarding the victim encoder or the type of backdoor trigger, as it activates in response to the occurrence of backdoor behavior. The contributions are summarized as follows:
  • We identify the pre-trained encoder mapping and its inverse’s capability of detecting potential backdoor behaviors from the objectives of backdoor attacks.
  • We propose DeDe to train a decoder as the inverse mapping of the encoder and reduce the reconstruction uncertainty by leveraging global embedding and constrained local embeddings.
  • Comprehensive experiments are conducted to demonstrate the effectiveness of our backdoor detection method in both CL and CLIP frameworks against various attack scenarios. In comparison to existing approaches, our method exhibits consistent performance and achieves high detection accuracy across all circumstances.
@@ 2.1 Pre-train Encoders via SSL
Self-supervised learning (SSL) plays a crucial role in training large upstream encoders as it frees massive dataset label annotations from human labor and potential human errors [36, 21, 16]. Various SSL techniques have been developed to facilitate real-world data mining, as the presence of unlabeled training instances is a common characteristic of large-scale datasets. It primarily seeks to leverage the inherent contextual relationships among the provided samples, which includes aspects such as spatial information, color information, the preservation of both local and global consistency, and other forms of prior knowledge to effectively perform a pre-training task. There are many different approaches to learning an encoder on an unlabeled dataset. Here, three widely recognized approaches are particularly noteworthy and relevant to our scope. Contrastive learning (CL) [5, 7, 37] exploits random data augmentations to transform unlabeled images into positive pairs and optimize their similarity. Masked auto-encoder (MAE) [19] is another promising approach that masks a portion of images and trains an encoder-decoder model to reconstruct the entire image. Recently, CLIP is an innovative form of CL using natural language supervision trained on a vast dataset consisting of 400 million image-text pairs sourced from the internet [31]. Without any explicit labels, it uses images and captions, such as a sentence describing the image content, to jointly train an image encoder and a text encoder. Its promising performance has drawn great attention into the field of multi-modal learning  [15]. Through these approaches, a pre-trained encoder is developed, and subsequently can be utilized in downstream tasks to map images to embeddings for downstream supervised tasks, such as classification.
@@ 2.2 Backdoor attacks and defences in SSL
Detecting backdoor in SSL is difficult due to the lack of data labels. There is no information available regarding which part of the data is triggered or the specific type of triggers present in the poisoned samples. This situation leads to a limited number of defensive strategies against SSL backdoor attacks. DECREE [13] is the first detection algorithm in the SSL settings. It requires the use of the shadow dataset, which contains backdoor samples, to optimize the trigger pattern. This is based on the observation that the distribution of clean samples and backdoor samples tends to cluster to different distributions. Subsequently, it determines whether the given encoder is implanted with a backdoor by assessing the proportion of the optimized trigger present in the entire image. ASSET [29] further proposes a detection method designed to identify poisoned samples in a given dataset across different learning paradigms, including SSL. It designs two opposite optimization problems over a small but clean dataset and the shadow dataset. By maximizing the optimization over the shadow dataset and minimizing over the clean dataset, the opposite optimization offsets the effect of the clean samples, leaving poisoned samples with relatively high loss, which provides a trace for the detection. [41] propose to identify triggered images by detecting outliers in the embeddings via the K-means algorithm and reverse-engineers the trigger in order to mitigate the backdoor’s effect. As for  [38], it presents an algorithm that distills a clean encoder from a compromised encoder through Min-Max optimization with the assumption of the backdoor’s existence.
CLIP [31] is frequently regarded as a contrastive learning paradigm for multi-modal tasks, and it has also been found susceptible to the backdoor threat. It is shown the backdoor may easily be injected into the CLIP model by merely $0.0001\%$ of poisoned samples  [4]. Sequentially, BadCLIP [26] proposes a novel attack that can bypass existing defenses and another BadCLIP [2] proposes a prompt-based backdoor attack to avoid malicious fine-tuning. In terms of countermeasures, there are two existing robust training frameworks; however, there are currently no detection methods effective in the presence of triggered samples within the dataset. As a robust training framework, CleanCLIP [3] designs an additional contrastive loss for each modal. ROCLIP [39] employs a candidate pool to match image embeddings with text embeddings at regular intervals, in which way, disrupting any potential associations between poisoned text-image pairs and realigning the correct relationships among text-image pairs. In BadCLIP [26], an active implanting algorithm still succeed in injecting the backdoor into the model.
@@ 3.1 What defines a good backdoor in encoders
To testify to the assertion and substantiate the challenge of stealthy backdoors, we provide a visualization of the embeddings associated with certain SSL backdoor attacks in Figure 1 via PCA. As shown, attacks like BadEncoder and CTRL (see Figure 4(a)) can be easily detected in the embedding space by any clustering algorithms, even for attacks imperceptible in the image space, such as CTRL. As the motivation suggested by [22, 35], an ideal encoder backdoor misleads the downstream tasks by altering the embedding of any triggered input to be close to the target-class representations. DRUPE is stealthier than its predecessors as visualized in Figure 1(c) and bypasses detection methods like DECREE. Therefore, we are motivated to develop a backdoor detector that is robust to any attacking scenarios, specifically for both out-of-distribution embeddings or ideally mixed embeddings.
@@ 3.2 Limitations in Existing Backdoor Defences
The techniques to defend against SSL backdoor attacks are still under-explored. It has been pointed out that the backdoor injected in contrastive learning is intrinsically different from its counterpart in supervised learning because the contrastive learning entangles the main task and the backdoor task, resulting in a mixture of inter-class and intra-class mapping [25]. DECREE as the first one to explore defending mechanism in SSL scenario, utilizes the observation on the embedding distribution that the backdoor samples are often clustered together and highly concentrated in a certain area, therefore, finding the small perturbation that can push a clean input towards the concentration area [12]. However, it suffers from two challenges. One is that it can only defend against attacks who demonstrate the observed pattern, which can be bypassed by new attack methods like  [35]. The other challenge is that DECREE is only applicable to patch-like triggers, leading to its ineffectiveness in defending attacks such as  [24, 33, 10]. Another defense method worth noting is ASSET[29], which uses the offset effect of two optimization tasks on the poisoned dataset and a manually chosen clean dataset to neutralize the clean samples influence, leaving the backdoor samples with a large loss to be detected. It has good performance in SSL attacks in general, however, it requires a sifted clean dataset from the training data, which might not always be available. It also shows a performance degeneration in more stealthy attacks and in different poison ratios. Moreover, there is a scarcity of effective backdoor countermeasures for SSL backdoor attacks. [25] has shown great challenges in terms of the learning dynamics and the detection scenarios. The existing works also have their own limitations. Hence, we are motivated to develop a defense mechanism to handle various backdoor attacks with stealthy embeddings or with imperceptible triggers from a more basic perspective of the attacker’s goal to manipulate the mappings of the encoder.
@@ 4.1 Attack and defense models
Attack scenario. We consider the two most practical scenarios from recent advancements of SSL paradigms.
On the model provider side, attackers may release poisoned samples to the web such that the automatically extracted data is poisoned similar to assumptions of  [33, 24, 27, 40] . SSL serves as a cheap and efficient way to train encoders without human censoring, leaving a loophole for attackers to pollute the collected dataset. In this case, the access to the meta-training dataset is granted, or at least a subset of it.
On the end user side, we consider someone who downloads pre-trained encoders to perform its own downstream tasks as an end user, which is also assumed by [22, 26]. However, when the third-party service provider may have injected a backdoor into the model without affecting the main utility, how can you know it as an end user of the pre-trained encoder without any knowledge of the training dynamics or training data? Hence, no training data is known but any publicly available dataset can serve as an auxiliary dataset.
Detection Goal. We assume that the attacker has already implemented the backdoor into the pre-trained encoder. How the backdoor is implemented and what type of trigger is are all agnostic. As a victim encoder needs to interact with a triggered sample to cause a backdoor behavior. We propose a detection method, DeDe, that detects the activation of backdoor behaviors when the victim model is inferencing a triggered input to a target embedding.
@@ 4.2 Key Idea
Figure 3: The Learning Goal of DeDe.
@@
An ideal backdoor attack should remain imperceptible at both the image level and the embedding level. At the image level, the attacker applies only a small perturbation (trigger) as a special distribution in image space, avoiding possible human inspections. At the embedding level, the attacker forces the victim encoder to learn the particular mapping from the trigger distribution to the targeted embedding. However by backdoor’s objective, it must force cross-class mapping from a class of image to the tatget-class embedding, which is illustrated as the red dot in Figure 3. Meanwhile, the backdoor’s secondary objective requires that the encoder’s utility upholds the main task performance, which suggests the compromised encoder still maps clean images to the respective embeddings as what a clean encoder does, as the blue dots in the figure. From the perspective of normal image-embedding mappings, it is from a sparse distribution in image space to dense clusters in the embedding space, which indicates that similar embeddings mean similar image features. If we find a reversed mapping from embedding clusters to images such that certain embedding regions map to certain visual features in the image, as the red arrows in the figure, the backdoor mapping is unlikely to be preserved. It is because, first, the embeddings space cannot encode as much information as image space to keep feature distribution and the backdoor distribution at the same time. Secondly, since the embedding distribution clusters, similar embeddings tend to map to similar image features. Since the dataset has a low poison rate, the learned mapping should be intra-class mapping as the majority does. If such an inverse is acquired, the backdoor embeddings map to the target-class images, leading to different visual features, which gives a clue to the backdoor detection by comparing the input and output images. For example, a cat image with the trigger is mapped to an embedding in the airplane’s region, if we find a point around the embedding and look for the point’s pre-image, it is highly probable that the corresponding image will depict an airplane.
@@
The idea of reversing the mapping is simple, however, finding such an inverse is challenging, because the inverted mapping is from a low-dimensional space to a high-dimensional space. The reconstruction is highly unpredictable since much information has degenerated by the encoder’s dimension reduction. Thus, how to reconstruct an image from the condensed embedding information is crucial. Inspired by MAE [19], we can allow some local information in image space to aid the reconstruction, which can give valuable reference for the object’s color, texture, and other local details. Therefore, we divide the reconstruction dependency into two parts, one is the global embedding and the other is the local patch embeddings. Fortunately, the triggered image is almost identical to the original image as the perturbation is small, moreover, the trigger is usually incomplete or neglected when we constrain the information from the image space by randomly selecting small patches. Leveraging the information from both global and local embeddings, we can train a decoder that reconstructs the image to approach the original one. Calling back to the above example, we reconstruct the pre-image from the embedding around the airplane region together with a few local patches of the original cat image. What would the reconstructed image be like? The decoder might reconstruct a cat-colored airplane or get confused and produce conflicting predictions. Then, when we compare the reconstruction with the actual pre-image, they are inconsistent and distinguishable by evaluating an error.

### v2

@@ Abstract
Self-supervised learning (SSL) is pervasively exploited in training high-quality upstream encoders with a large amount of unlabeled data. However, it is found to be susceptible to backdoor attacks merely via polluting a small portion of training data. The victim encoders associate triggered inputs with target embeddings, e.g., mapping a triggered cat image to an airplane embedding, such that the downstream tasks inherit unintended behaviors when the trigger is activated. Emerging backdoor attacks have shown great threats across different SSL paradigms such as contrastive learning and CLIP, yet limited research is devoted to defending against such attacks, and existing defenses fall short in detecting advanced stealthy backdoors. To address the limitations, we propose a novel detection mechanism, DeDe, which detects the activation of backdoor mappings caused by triggered inputs on victim encoders. Specifically, DeDe trains a decoder for any given SSL encoder using an auxiliary dataset (which can be out-of-distribution or even slightly poisoned), so that for any triggered input that misleads the encoder into the target embedding, the decoder generates an output image significantly different from the input. DeDe leverages the discrepancy between the input and the decoded output to identify potential backdoor misbehavior during inference. We empirically evaluate DeDe on both contrastive learning and CLIP models against various types of backdoor attacks. Our results demonstrate promising detection effectiveness over various advanced attacks and superior performance compared over state-of-the-art detection methods.
@@ 1 Introduction
In this paper, we propose DeDe, an effective detector for detecting backdoor samples in SSL settings. To address the limitations of existing detection methods and the emergence of increasingly advanced attacks, the proposed approach directly leverages the relationships between embeddings and images. DeDe aims to establish a region-to-region mapping to invert the normal mapping of an encoder, thereby enabling the identification of malicious behaviors that deviate from the normal mapping. Encoders’ normal utility can be considered as the correct mapping from similar data (usually from the same class) to a certain region of the embedding space [37]. Embeddings of similar images are observed to converge to a clustered region in the vector space; conversely, backdoor encoders aim to map triggered images to a different embedding region instead. Our approach exploits this property to find a bijective mapping between classes of images and classes of embeddings, where normal utility leads to intra-class mapping, while backdoor behaviors introduce inter-class mapping.
Overall, the proposed detection approach offers three significant advantages: 1) the proposed approach is effective and robust against various advanced backdoor attacks, including which use the invisible trigger or stealthy embedding; 2) it does not rely on the availability of a clean dataset or an in-distribution dataset; and 3) it is a non-invasive detection method that requires no prior information regarding the victim encoder or the type of backdoor trigger, as it solely alerts in response to the occurrence of backdoor behavior. The contributions are summarized as follows:
  • We identify the pre-trained encoder mapping and its inverse’s capability of detecting potential backdoor behaviors based on the intrinsic objectives of backdoor attacks.
  • We propose DeDe as a novel backdoor detection mechanism utilizing the inverse mapping of an encoder. It trains a decoder for the inverse mapping of the encoder and reduces the reconstruction uncertainty by leveraging global embedding and constrained local image features.
  • Comprehensive experiments are conducted to demonstrate the effectiveness of our backdoor detection method in both CL and CLIP frameworks against various attack scenarios. In comparison to existing defense approaches, our method exhibits consistent performance and achieves high detection accuracy across all circumstances.
@@ 2.1 Pre-train Encoders via SSL
Self-supervised learning (SSL) plays a crucial role in training large upstream encoders as it frees massive dataset label annotations from human labor and potential human errors [37, 22, 16]. Various SSL techniques have been developed to facilitate real-world data mining, as the presence of unlabeled training instances is a common characteristic of large-scale datasets. It primarily seeks to leverage the inherent contextual relationships among the provided samples, which includes aspects such as spatial information, color information, the preservation of both local and global consistency, and other forms of prior knowledge to effectively perform a pre-training task. There are many different approaches to learning an encoder on an unlabeled dataset. Here, three widely recognized approaches are particularly noteworthy and relevant to our scope. Contrastive learning (CL) [5, 7, 38] exploits random data augmentations to transform unlabeled images into positive pairs and optimize their similarity. Masked auto-encoder (MAE) [19] is another promising approach that masks a portion of images and trains an encoder-decoder model to reconstruct the entire image. Recently, CLIP is an innovative form of CL using natural language supervision trained on a vast dataset consisting of 400 million image-text pairs sourced from the internet [32]. Without any explicit labels, it uses images and captions, such as a sentence describing the image content, to jointly train an image encoder and a text encoder. Its promising performance has drawn great attention into the field of multi-modal learning  [15]. Through these approaches, a pre-trained encoder is developed, and subsequently can be utilized for downstream supervised tasks, such as classification.
@@ 2.2 Backdoor attacks and defences in SSL
Detecting backdoor in SSL is difficult due to the lack of data labels. There is no information available regarding which part of the data is triggered or the specific type of triggers present in the poisoned samples. This situation leads to a limited number of defensive strategies against SSL backdoor attacks. DECREE [13] is the first detection algorithm in the SSL settings. It requires the use of the shadow dataset, which contains backdoor samples, to optimize the trigger pattern. This is based on the observation that the distribution of clean samples and backdoor samples tends to cluster to different distributions. Subsequently, it determines whether the given encoder is implanted with a backdoor by assessing the proportion of the optimized trigger present in the entire image. ASSET [30] further proposes a detection method designed to identify poisoned samples in a given dataset across different learning paradigms, including SSL. It designs two opposite optimization problems over a small but clean dataset and the shadow dataset. By maximizing the optimization over the shadow dataset and minimizing over the clean dataset, the opposite optimization offsets the effect of the clean samples, leaving poisoned samples with relatively high loss, which provides a trace for the detection. It is worth noting that the pipeline of backdoor sample detection like ASSET is similar to a few works in out-of-distribution (OoD) detection [21, 40, 41], though with different objectives. [44] propose to identify triggered images by detecting outliers in the embeddings via the K-means algorithm and reverse-engineers the trigger in order to mitigate the backdoor’s effect. As for  [39], it presents an algorithm that distills a clean encoder from a compromised encoder through Min-Max optimization with the assumption of the backdoor’s existence.
CLIP [32] is frequently regarded as a contrastive learning paradigm for multi-modal tasks, and it has also been found susceptible to the backdoor threat. It is shown the backdoor may easily be injected into the CLIP model by merely $0.0001\%$ of poisoned samples  [4]. Sequentially, BadCLIP [27] proposes a novel attack that can bypass existing defenses and another BadCLIP [2] proposes a prompt-based backdoor attack to avoid malicious fine-tuning. In terms of countermeasures, there are two existing robust training frameworks; however, there are currently no detection methods effective in the presence of triggered samples within the dataset. As a robust training framework, CleanCLIP [3] designs an additional contrastive loss for each modal. ROCLIP [42] employs a candidate pool to match image embeddings with text embeddings at regular intervals, in which way, disrupting any potential associations between poisoned text-image pairs and realigning the correct relationships among text-image pairs. In BadCLIP [27], an active implanting algorithm still succeeds in injecting the backdoor into the model.
@@ 3.1 What defines a good backdoor in encoders
To testify to the assertion and substantiate the challenge of stealthy backdoors, we provide a visualization of the embeddings associated with certain SSL backdoor attacks in Figure 1 via PCA. As shown, attacks like BadEncoder and CTRL (see Figure 4(a)) can be easily detected in the embedding space by any clustering algorithms, even for attacks imperceptible in the image space like CTRL. As the motivation suggested by [23, 36], an ideal backdoor in encoder misleads the downstream tasks by altering the embedding of triggered input to be close to the target-class representations. DRUPE is stealthier than its predecessors as visualized in Figure 1(b), 1(c) and bypasses detection methods like DECREE. Therefore, we are motivated to develop a backdoor detector that is robust to any attacking scenarios, specifically for either out-of-distribution embeddings or ideally mixed embeddings.
@@ 3.2 Limitations in Existing Backdoor Defences
The techniques to defend against SSL backdoor attacks are still under-explored. It has been pointed out that the backdoor injected in contrastive learning is intrinsically different from its counterpart in supervised learning because the contrastive learning entangles the main task and the backdoor task, resulting in a mixture of inter-class and intra-class mapping [26]. DECREE as the first one to explore the defending mechanism in SSL scenario, utilizes the observation on the embedding distribution that the backdoor samples are often clustered together and highly concentrated in a certain area, therefore, finding the small perturbation that can push a clean input towards the concentration area [12]. However, it suffers from two challenges. One is that it can only defend against attacks who demonstrate observable embedding patterns, which can be bypassed by newly developed imperceptible attacks like  [36]. The other challenge is that DECREE is only applicable to patch-like triggers, leading to its ineffectiveness in defending attacks such as  [25, 34, 10]. Another defense method worth noting is ASSET[30], which uses the offset effect of two optimization tasks on the poisoned dataset and a manually chosen clean dataset to neutralize the clean samples influence, leaving the backdoor samples with a large loss to be detected. It has good performance in SSL attacks in general, however, it requires a sifted clean dataset from the training data, which might not always be available. It also shows a performance degeneration in more stealthy attacks and in different poison ratios. Moreover, there is a scarcity of effective backdoor countermeasures for SSL backdoor attacks. [26] has shown great challenges in terms of the learning dynamics and the detection scenarios. The existing works also have their own limitations. Hence, we are motivated to develop a defense mechanism to handle various backdoor attacks with stealthy embeddings or with imperceptible triggers from a more basic perspective of the attacker’s goal to manipulate the mappings of the encoder.
@@ 4.1 Attack and defense models
Attack scenario. We consider the two most practical scenarios from recent advancements in SSL paradigms.
On the model provider side, attackers may release poisoned samples to the web such that the automatically extracted data is poisoned similar to assumptions of  [34, 25, 28, 43] . SSL serves as a cheap and efficient way to train encoders without human censoring, leaving a loophole for attackers to pollute the collected dataset. Hence, access to the meta-training dataset is granted, or at least a subset of it.
On the end user side, consider someone who downloads pre-trained encoders to perform its own downstream tasks as an end user, which is also assumed by [23, 27]. However, when the third-party service provider may have injected a backdoor into the model without affecting the main utility, how can you, as an end user, know it without any knowledge of the training dynamics or training data? Hence, no training data is known but any publicly available dataset can serve as an auxiliary dataset.
Detection Goal. We assume that the attacker has already implemented the backdoor into the pre-trained encoder. How the backdoor is implemented and what type of trigger is both are agnostic. As a victim encoder needs to interact with a triggered sample to cause a backdoor behavior, we propose a detection method, DeDe, that detects the activation of backdoor behaviors when the victim model is inferencing a triggered input to a target embedding.
@@ 4.2 Key Idea
Figure 3: The Learning Goal for Decoder of DeDe.
@@
In essence, SSL spontaneously maps images with similar features to clusters, which is determined by its training process. We present t-SNE plots for different SSL paradigms in the Appendix. However, backdoor mappings force cross-class mapping from a class of image to the target-class embedding, which is illustrated by the red dots in Figure 3. As required by backdoor’s objective, an ideal backdoor should remain imperceptible at both the image level and the embedding level. At the image level, the attacker applies only a small perturbation (trigger) as a special distribution in image space, avoiding possible human inspections. At the embedding level, the attacker forces the victim encoder to learn the particular mapping for triggered samples to the targeted embedding while keeping clean images to the respective embeddings as the blue dots in Figure 3. A naive instinct is that embeddings from the same cluster should represent similar visual features in images. If we learn the reversed mapping from embedding clusters to images as the red arrows in the figure, the reversed mapping should project certain embedding regions to certain visual features in the image, then the backdoor mapping is unlikely to be preserved. It is because, first, the image space is much more sparse than the embedding space, so condensed embedding is unlikely to preserve a backdoor distribution in image. Secondly, since the embedding distribution forms clusters, similar embeddings tend to map to similar image features, but the backdoor feature is usually imperceptible and dissimilar to image features. For example, a cat image with the trigger is maliciously mapped to an embedding in the airplane’s region, if we find a point around its embedding and look for the point’s pre-image, it is highly probable that the corresponding image will depict an airplane.
@@
The idea of reversing the mapping is simple, however, finding such an inverse is challenging, because the inverted mapping is from a low-dimensional space to a high-dimensional space. The reconstruction is highly unpredictable since much information has degenerated by the encoder’s dimension reduction. Thus, how to reconstruct an image from the condensed embedding information is crucial. Inspired by MAE [19], we can allow some local information in image space to aid the reconstruction, which can give valuable reference for the object’s color, texture, background and other local details. More critically, a triggered image is almost identical to the original image as the perturbation has to be small, as a result, the trigger is usually incomplete or neglected when we constrain limited information from the image space by randomly selecting small patches. Therefore, we divide the reconstruction dependency into two parts, one is the global embedding and the other is the local patches. Leveraging the information from both global and local embeddings, we can train a decoder that reconstructs the image to approach the original one. Calling back to the above example, we reconstruct the pre-image from the embedding around the airplane region together with a few local patches of the original cat image. What would the reconstructed image be like? The decoder might reconstruct a cat-colored airplane or get confused to produce nonsensical prediction. In either way, when we compare the reconstruction with the actual pre-image, they are inconsistent and distinguishable by simply evaluating an error. We present the workflow of DeDe in Figure 2.